OpenStudy (anonymous):

Simplify the following: 3xy^4z^5 __________ 30x^2y^4z^4

OpenStudy (luigi0210):

\[\Huge \frac{a^x}{a^y}=a^{x-y}\]

OpenStudy (luigi0210):

So separate them: \[\LARGE \frac{3}{30}*\frac{x}{x^2}*\frac{y^4}{y^4}*\frac{z^5}{z^4}\]

OpenStudy (anonymous):

the top xy^4 is on term the bottom is two term x^2 and y^4 so separating the the top xy^4 won't work

OpenStudy (luigi0210):

Yea it would, it's still multiplication. :)